OpenStudy (anonymous):

18x^2 y + 24x^3 y^4 - 12x^7 y^2 OVER 6x^2 y

OpenStudy (anonymous):

Uhh, it's just an expression.

OpenStudy (anonymous):

Do you want to know how to simplify it?

OpenStudy (anonymous):

\[{18x^2 y + 24x^3 y^4 - 12x^7 y^2 \over 6x^2 y}={18x^2 y \over 6x^2 y}+{24x^3 y^4 \over 6x^2 y}-{12x^7 y^2 \over 6x^2 y}\]Use the property\[\huge {x^a \over x^b}=x^{a-b}\]Simplify the constants just like back in grade school.

OpenStudy (anonymous):

yeah I need an explation to simplify it

OpenStudy (anonymous):

well every term in the numerator has a factor of 6, every term in the numerator has a factor of x^2, and every term in the numerator has a factor of y. so basically you can factor out the something equal to the denominator from the numerator and cancel it

OpenStudy (anonymous):

$$\frac{18x^2y+24x^3y^4-12x^7y^2}{6x^2y}$$Our first step is to notice that every term in the top has a \(6x^2y\) in common, so let's factor it out:$$\frac{6x^2y[3+4xy^3-2x^5y]}{6x^2y}$$Now we can cancel out \(6x^2y\) as its a common factor in both the numerator and denominator:$$3+4xy^3-2x^5y$$